OpenStudy (anonymous):

Human hair grows at a rate of approximately three-twentieths centimeters per year. How many centimeters does it grow in two-thirds of a year? one-tenth centimeters forty-nine sixtieths centimeters nine-fortieths centimeters thirty-one sixtieths centimeters

OpenStudy (anonymous):

@phi can you help me please if you can

OpenStudy (phi):

can you write a ratio ?

OpenStudy (anonymous):

@phi well I have trouble with ratios

OpenStudy (phi):

another way is use rate * time = distance (or "growth" in this case) what is the rate ? what is the time ?

OpenStudy (anonymous):

rate: \[\frac{ 3 }{ 20 }\] time: \[\frac{ 2 }{ 3 }\]

OpenStudy (phi):

yes, but you should put in the units. now multiply rate * time \[ \frac{3 \text{ cm}}{20\cancel{\text{ yr}}} \cdot \frac{2}{3} \cancel{\text{yr}}= ?\]

OpenStudy (anonymous):

@phi ok.. so what do I do..?

OpenStudy (phi):

multiply rate * time

OpenStudy (anonymous):

ok hold on a second

OpenStudy (anonymous):

it gives me 0.1 which is 1/10

OpenStudy (anonymous):

and 1/10 is one of my answer choices so that the answer right?

OpenStudy (phi):

yes. when you multiply fractions, you multiply top times top and bottom times bottom \[ \frac{3 \cdot 2}{20 \cdot 3 } \] notice you can divide top and bottom by 3 (it's easier to simplify first) you get \[\frac{2}{20} \] now notice you can divide top and bottom by 2 to simplify more. you you get \[ \frac{1}{10} \text{ or } 0.1 \]

OpenStudy (phi):

and the answer is 0.1 cm or 1/10 cm

OpenStudy (anonymous):

ty